what is the easiest way to do this problem?Problem 11 < 11 of 14 C Review Constants Periodic Table Three identical very dense masses of 3600 kg each are placed on the x
Answer #1

net force exert on the mass placed at origin

F = force exerted by mass at x = - 120 cm - force exerted by mass at x = 450 cm

F = Gm ( M1/r1^2 - M2 / r2^2)

F = 6.67* 10^-11* 3600* ( 3600 / 1.2^2 - 3600 /4.5^2)

F = 5.576* 10^-4 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck